OpenStudy (anonymous):

twenty two slips of paper are each marked with a different letter of the alphabet and placed in a basket a slip is pulled out its letter recorded and the slip is replaced this is done 4 times find the probability that a word with no repeptition of letters is formed

OpenStudy (anonymous):

did you get the last one?

OpenStudy (anonymous):

ya i did. i need help with this one!

OpenStudy (anonymous):

the last one was much harder

OpenStudy (anonymous):

Help me with this one

OpenStudy (anonymous):

they have to be all different, so it is going to be \[\frac{20\times 19\times 18\times 17}{20^4}\]

OpenStudy (anonymous):

that is a mistake, there are 22 letters, not 20!

OpenStudy (anonymous):

\[\frac{22\times 21\times 20\times 19}{21^4}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Hmmmmmmm... not sure I believe...